Your car’s catalytic converter loses 50 ng Pd for each km you drive. How many miles do you drive before you lose 1 g?
1
Expert's answer
2019-04-12T07:49:28-0400

m(Pd) = 50 ng = 50×10-9 g = 5×10-8 g.


5×10-8 g – 1 km

1 g          – X km

X = 1/5×10-8 km = 2×107 km.


1 mile   – 1.609 km

Y miles – 2×107 km

Y = 2×107/1.609 miles = 1.243×107 miles.